### v1

# Context Matters: Peer-Aware Student Behavioral Engagement Measurement via VLM Action Parsing and LLM Sequence Classification

Abstract
Understanding student behavior in the classroom is essential to improve both pedagogical quality and student engagement. Existing methods for predicting student engagement typically require substantial annotated data to model the diversity of student behaviors, yet privacy concerns often restrict researchers to their own proprietary datasets. Moreover, the classroom context, represented in peers’ actions, is ignored. To address the aforementioned limitation, we propose a novel three-stage framework for video-based student engagement measurement. First, we explore the few-shot adaptation of the vision-language model for student action recognition, which is fine-tuned to distinguish among action categories with a few training samples. Second, to handle continuous and unpredictable student actions, we utilize the sliding temporal window technique to divide each student’s 2-minute-long video into non-overlapping segments. Each segment is assigned an action category via the fine-tuned VLM model, generating a sequence of action predictions. Finally, we leverage the large language model to classify this entire sequence of actions, together with the classroom context, as belonging to an engaged or disengaged student. The experimental results demonstrate the effectiveness of the proposed approach in identifying student engagement.
1 Introduction
Students’ active participation and engagement in the classroom are crucial for successful learning outcomes [13]. Student engagement is a multidimensional construct that encompasses behavioral, emotional and cognitive engagement [17]. Behavioral engagement refers to students’ self-directed actions to get access to the curriculum, such as studying and doing homework, hand and body movements while attending lectures, as well as engaging cooperatively in classroom activities [4]. Emotional engagement, on the other hand, refers to students’ feelings about their learning experience, teachers, and peers [27]. This involves feelings of happiness in learning, disinterest in the subject matter, or frustration and struggle to understand. Lastly, cognitive engagement is the psychological investment in academic achievement, which is usually revealed by a student’s ability to comprehend, organize, and analyze in the act of deep learning to fully understand the material [10].
Behavioral engagement is a key indicator of student participation and attentiveness, with a direct impact on academic performance [5]. During a lecture, a student’s behavior can reveal whether they are on- or off-task. While a disengaged student engages in off-task activities like playing with a mobile device and checking the time, an engaged student typically listens, takes notes, and types on a laptop. Traditional assessments of this engagement depend on manual classroom observations [28, 19]. While informative, these observations are labor-intensive and subject to observer bias, which limits their scalability in large educational settings. A key aspect of engagement annotation is its temporal resolution, which defines the time scale at which engagement is evaluated. Following the standard classroom observation protocols [28, 19], we adopt a 2-minute time resolution, i.e., the length of each video sample is set to match a 2-minute time window.
Automating student behavioral engagement in classrooms has received considerable attention due to advances in computer vision and deep learning. Existing studies for student behavior analysis can be categorized according to the target outcome into action recognition, learning analytics, and engagement measurement. Methods for recognizing student actions focus on identifying a particular group of students’ behaviors without conducting additional analysis, providing instructors with minimal useful information. Furthermore, a considerable number of studies [20, 36] formulated student behavior analysis as single- or multi-object detection, ignoring the complexity and diversity of student behaviors. Learning analytics approaches, which use current computer vision techniques to analyze student behavior, provide a set of visual features but little useful information for instructors. In these approaches, a specific student’s actions, gaze, and/or facial expression are usually employed in estimating the student’s engagement.
Unlike traditional Human Action Recognition (HAR) models, which were trained on clips that have been trimmed to include only one action, detecting student actions over arbitrary time intervals necessitates the analysis of continuous, untrimmed video streams in which the execution and order of actions are critical. To bridge this gap, research has shifted toward methods for temporal action localization (TAL) [29, 34], which identify and locate the start and end times of sparse, intermittent actions within long, untrimmed videos. However, TAL tends to ignore “background” segments (e.g., listening, looking to the side/back), which are important in engagement analysis. Moreover, Yu et al., [34] leveraged LLM to generate a behavior analysis report without estimating students’ engagement. In order to measure students’ behavioral engagement, Abdelkawy et al. [2] created a histogram of the students’ actions over an arbitrary time period. The video clip was divided into overlapping segments using the sliding temporal window technique over a 2-minute period. These segments are then fed into a 3D CNN model for action recognition. A classifier that assesses the student’s level of engagement uses both this histogram and the student’s gaze as inputs. However, the student’s actions throughout such a time interval are not manually annotated. Thus, the evaluation of the behavioral engagement classifier remains incomplete.
Figure 1: Context matters in engagement classification. Under traditional approaches, a student’s action sequence showing continuous typing on a laptop may be labeled as engaged. However, our peer-aware model classifies the same sequence as disengaged when surrounding students are not engaged in similar note-taking behavior.
The limitations of the three types of approaches are:
  • There is a scarcity in publicly available datasets for students actions and engagement, since in each studey, the authors collect their proprietary dataset due to privacy concerns.
  • The context, which is defined as actions performed by peers while the student of interest is performing a certain action, is underestimated or ignored (see Fig. 1).
Since student engagement depends not only on their actions but also on the current classroom activity and the behavior of their peers, classroom context is crucial to accurately assessing student engagement. This is especially important as contemporary pedagogy increasingly moves beyond passive lecture-based learning. ICAP framework [10] characterizes learning activities along a Passive–Active–Constructive–Interactive hierarchy, where interactive tasks generally reflect deeper cognitive engagement. The same observable behavior can therefore signal opposite engagement states depending on context. A student turning to talk to a peer during a traditional lecture, for instance, probably shows disengagement, but doing the same during a peer-discussion activity that is prompted shows engagement. When interpreting any student action in a classroom, context is important. For example, when a student asks a question, other students are expected to turn to listen to their peer; when a student is asked to work independently on a problem or possibly with a partner, they are engaged by focusing intently on their laptop for problem solving; when the instructor is lecturing or demonstrating a problem solution, on the other hand, they are disengaged by focusing on their laptop while accessing their social media feeds.
We propose a novel three-stage framework for classifying behavioral engagement based on student body movement in order to overcome the aforementioned drawbacks. In order to classify human actions in trimmed videos with a limited number of labeled samples per category, we first perform few-shot adaptation for the vision-language model [23]. Second, because the student actions are continuous, we divide each 2-minute time-resolution student video into non-overlapping segments using the sliding temporal window technique. The VLM model (from stage 1) is then used to assign a predefined label to each segment, resulting in a time-ordered sequence of predicted actions. Third, an LLM uses this sequence of actions and the classroom context as input to determine whether a student is engaged or disengaged.
To the best of our knowledge, prior studies have not investigated whether current vision-language models (VLMs) and large language models (LLMs) can effectively recognize the student’s class actions and measure behavioral engagement based on identified actions, respectively. Accordingly, this paper addresses the research questions:
  1. 1.
    Compared to traditional action classification, is it harder for VLMs to classify actions of students in-class?
  2. 2.
    Which action sequence or action histogram most accurately depicts students’ behavior?
Our contributions are two-fold:
  • Develop a novel framework for measuring student behavioral engagement, which integrates VLM for action recognition and LLM for engagement classification. To address the lack of publicly available datasets for student actions and engagement, we perform a few-shot tuning for VLM to improve student action recognition, and we use the LLM in a zero-shot prompting for engagement classification.
  • Create a novel student behavior dataset with two components: 1) trimmed videos (13 categories, controlled environment) with annotated students’ actions; and 2) untrimmed videos (3 lectures, 11 students) with annotated students’ engagement. The experts’ annotations of this dataset include ”engaged/disengaged” labels for each 2-minute video, and dense temporal action labels for each segment within them.
2 Related Work
2.1 Measuring Student Engagement
The emerging and promising real-time engagement measurement approach, enabled by advances in computer vision and machine learning, allows tracking a student’s engagement at fine-grained temporal resolutions on the order of seconds or minutes through discrete, objective indicators. Existing research of student behavior analysis in classrooms can be grouped into three categories: student action recognition, learning analytics, and engagement measurement. While student action recognition approaches focus on the student’s actions from trimmed video or a single image, learning analytics approaches rely on facial expressions, student actions, and/or head gaze. For example, Sheng et al. [25] proposed an enhanced YOLOv8-based detector for localizing and classifying student behaviors from single classroom images. Their framework incorporates a Multi-Scale Large Kernel Convolution Module (MLKCM) to extract multiscale spatial features and a Progressive Feature Optimization Module (PFOM) that segments the channel dimension of the input feature map for improved feature refinement. Ahuja et al. [3] presented a classroom sensing system (EduSense) that provides a set of features, including detection of hand raise, hand on face, cross arms, smile, head orientation, and speech duration for both the teacher and the students. The authors employed students’ skeleton joints and students’ facial landmarks to extract these visual features due to the presence of much visual noise, occlusion in the classroom, and privacy preservation. The student’s actions are recognized frame-wise, which leads to a high false-positive rate in recognition, especially for hand raising. Line et al.[21] recognized student behavior in the classroom based on upper body skeleton joints. Their framework includes three steps: Extracting students’ poses using bottom-up pose estimation (Openpose) and removing the incorrect connection of students’ joints using YoloV4 object detection, generating feature vector using normalized joint locations, joint distances, and bone angles to represent student posture, and using four fully connected layer neural network to classify the student feature vector into four classes. The main limitations of their framework are that it recognizes student actions from single images rather than video segments, and it considers actions with different engagement levels, such as writing and playing with a phone as one class (bowing).
Student engagement studies, which employ student actions and/or head gaze to measure the student behavioral engagement, provide the teachers with actionable insights. In [35], the authors proposed a framework that processes classroom videos to first estimate each student’s 2D pose per frame and then apply a Temporal Action Detection (TAD) [16] module to produce a temporal sequence of actions for each student. These per-student action sequences are subsequently fed into an LLM to generate behavioral reports. The TAD module was trained on the NTU RGB+D dataset [24], using only action classes relevant to classroom contexts. However, this approach has two main limitations. First, it assumes that each student’s behavior is independent of others, disregarding the broader classroom context. Second, the model does not predict explicit engagement labels; instead, it outputs textual behavioral reports that still require human interpretation to extract meaningful insights. Abdelkawy el al.[2] utilized a random forest classifier to detect whether the student is on-task or off-task based on a histogram of actions, together with head gaze. Each student’s 2-minute-long video was divided into overlapping segments using the sliding temporal window technique; the student skeleton sequence of each segment is assigned a predefined action label using 3D CNN model and then the histogram is created from these identified actions. Nevertheless, the student’s actions during the given period are not manually annotated. Consequently, the evaluation of the behavioral engagement classifier is still incomplete. Unlike these studies, we analyze the entire 2-minute video to get the student’s sequence of actions, which are used for engagement measurement. Moreover, we constructed a dataset that has dual expert annotations: engaged/disengaged for each 2-minute video and its corresponding frame-wise action labels to measure the performance of the engagement classifier. We also investigated the few-shot adaptation of the vision-language model for student action recognition, which learns to differentiate between action classes with a few training samples to overcome the challenge of the scarcity of annotated data required to model the diversity of student behaviors.
Figure 2: Overview of the proposed framework. Students’ tubelets are extracted from a two-minute video, and each student’s tube is divided into non-overlapping segments using a sliding temporal window technique. The action sequence is then produced by the fine-tuned VLM model. LLM uses the generated action sequence to classify engagement after integrating it with peers’ action sequences.
2.2 Video Action Recognition
Action recognition aims to assign a predefined action label to a trimmed video, which contains a single action. The majority of existing action recognition approaches are based on RGB video streams. Existing RGB video-based methods can be categorized into four classes based on the network architecture: two-stream 2D convolutional neural network (CNN), 3D CNN, compute-efficient, and transformer-based methods [1]. Two-stream 2D CNN-based methods [26, 32] learn the appearance and motion features separately using two 2D CNN streams. 3D CNN-based methods [9, 15] are introduced to learn spatiotemporal features by temporally extending the convolution and pooling kernels of a 2D CNN. However, 3D CNNs are computationally intensive. To achieve the computation/accuracy tradeoff, compute-efficient networks [31, 14] are introduced. Transformer-based methods [7, 6] adapt the de facto vision transformer to model the long-term temporal dynamic in a video sequence. While most of these methods are unimodal, ignoring the semantic information of action categories, works such as ActionCLIP [33] and XCLIP [22] adapt the CLIP to vidoe action recognition through fusing the vision-language representation of CLIP with introduced temporal modeling components. Rasheed et al.[23] found that these introduced components can negatively impact CLIP’s generalization capability, whereas full fine-tuning of CLIP on the Kinetics-400 dataset yields competitive performance. In this work, we adapt the work of [23] in a few-shot setting to boost the performance of student classroom action recognition.
2.3 Large Language Models
Large language models (LLMs) are Transformer-based models with hundreds of billions of parameters, trained on large-scale text corpora (e.g., GPT-3 [8], LLaMA [12], Gemma [30], etc.). They exhibit generalization abilities, enabling them to perform tasks in zero-shot settings without requiring task-specific training examples. While several works employed LLMs in education for multiple-choice question generation and assignment grading, there is less work that leverages the reasoning capability of LLMs in student behavior analysis. Yu et al. [34] utilized an LLM to generate a pedagogical report based on the structured temporal action information for each student, which includes action categories and their timestamps. In this work, we use the LLM to assign an engagement level to each student’s sequence of actions.
3 Method
3.1 Overview of the proposed method
The proposed framework consists of three main steps: segment-wise action recognition, VLM action parsing, and engagement classification (Fig. 2). First, a vision-language model is adapted in a few-shot manner using a small set of labeled trimmed student action clips to improve recognition on continuous student videos. Second, the student’s untrimmed video is split into non-overlapping segments using a sliding-window technique, and the fine-tuned VLM assigns an action label to each segment. Consecutive identical predictions are merged to form an ordered action sequence that preserves both action identity and duration. Finally, this action sequence, along with the classroom context, is provided to an LLM to determine whether the student is engaged or disengaged. Representing student’s behavior during a 2-minute interval as a temporal sequence rather than a histogram, where the duration of each action is accumulated, allows the model to capture not only how long each action occurs but also their ordering, which is shown to be crucial for interpreting engagement. For example, a student may still be deemed engaged if they take notes for the majority of the time period but play on their phone for one continuous 25-second block (e.g., at the beginning or end). However, a student is considered disengaged if they alternate between taking notes and making frequent quick phone checks because this pattern of frequent disruption shows an inability to maintain cognitive focus on the task. Therefore, the temporal sequence and context of actions more accurately reflect engagement than the total amount of time spent in each action (e.g., via a histogram).
Furthermore, when representing student action as a sequence, together with peers’ action sequences, the LLM predictions are easier to be interpreted. Fig.3 shows examples of students’ actions aligned with classroom actions for engaged (a) and disengaged (b, c) students.
Figure 3: a) for engaged student, b, c) for disengaged student
3.2 Actions dictionary
Since this study is conducted on college students in introductory engineering STEM classes, the actions they perform during classroom time, from which student behavior engagement can be estimated, are limited. We adopted classroom observation protocols in [2], and we came up with the following 13 actions:
  • Eating meal/snack,
  • Writing on notebook/tablet,
  • Typing on a laptop,
  • Playing with mobile phone,
  • Looking to the side/back,
  • Looking down w/o reading/writing,
  • Looking at laptop screen (not typing).
  • Raising hand,
  • Rubbing face,
  • Reading,
  • Drinking,
  • Yawning,
  • Listening,
Fig. 4 shows samples of such actions.
Figure 4: Samples of students’ actions during class time.
3.3 Student action recognition
We investigated few-shot learning for action recognition using vision-language models (VLMs). In particular, we use Video Fine-tuned CLIP (ViFi-CLIP) [23] because of its efficiency and simplicity. The ViFi-CLIP adapted CLIP to the video domain through fine-tuning both the CLIP image and text encoders. For a video segment $S_{i}\in\mathbb{R}^{TxHxWxC}$, the embeddings of its T frames are obtained using the CLIP image encoder and then average-pooled to get the video-level representation $v_{i}\in\mathbb{R}^{D}$ . The corresponding class label $C_{i}$ is embedded in a prompt (e.g., ‘a photo of a <<category>>’) and passed through the CLIP text encoder to produce a text embedding $t\in\mathbb{R}^{D}$. We fine-tuned the ViFi-CLIP model on the student actions dataset in a few-shot setting with 16 video clips for each action category by maximizing the cosine similarity between a batch of video-level embeddings $v_{i}$ and their corresponding text embeddings $t_{i}$ using Cross Entropy loss with entropy regularization,
$L_{total}=-\frac{1}{N}\sum_{i}^{N}log(p_{i})+\frac{1}{N}\sum_{i}^{N}(-\sum_{k=1}^{C}p_{i,k}log(p_{i,k}))$ (1)
, where
$p_{i}=\frac{exp(sim(v_{i},t_{i})/\tau}{\sum_{j}exp(sim(v_{j},t_{j}/\tau)}$ (2)
3.4 Student engagement classification
To determine whether a student is engaged or disengaged, we leverage the LLM as a classifier in a zero-shot prompting. Following the prompt-based learning paradigm, the LLM is conditioned on a prompt $x_{prompt}$ to generate a predefined textual label $y\in Y$ (e.g., engaged, disengaged). The prompt $x_{prompt}$ is constructed by concatenating a task description $x_{desc}$ and input $x_{input}$ as $x_{desc}$$x_{input}$}. An example of the task description $x_{desc}$, which describes the engagement classification based on the student’s actions and classroom context, is shown in Fig. 6 An example of the input $x_{input}$, which includes the student action sequence with the duration of each action, along with the classroom context, represented in the majority of peers’ actions during the time interval, can be represented as:
  • Student actions: { writing on tablet (00:00-00:20); listening (00:20-01:05; etc.}
  • Classroom context: {listening (00:00-00:10); writing (00:10-00:50); etc.}
Figure 6: An example of the task description $x_{desc}$
4 Experiments
4.1 Experimental Setup
Benchmarks. The lack of a public benchmark for classroom engagement analysis makes it necessary to create one. So, we created a behavioral engagement dataset with two subsets of annotated videos—one for student actions and the other for student engagement—in order to independently assess each step of the proposed framework. All videos were recorded from a wall-mounted 4K camera at 15 fps in classrooms of 6–11 students. For the student action dataset, we recorded 7 subjects performing 13 classroom-related actions in a controlled environment. This resulted in 208 trimmed clips for training (16 per action) and 47 trimmed video clips for testing across the 13 action categories. This trimmed video dataset is designed to evaluate the proposed action recognition stage (Stage 1). The student engagement dataset consists of videos from three half-hour lectures involving eleven students. Experts in psychology and education labeled each two-minute clip as either engaged or disengaged. Each two-minute video is also temporally partitioned, and every resulting segment is assigned a predefined action label. This untrimmed video dataset is used to evaluate the remaining stages: proposed temporal segmentation (Stage 2) and engagement classification (Stage 3).
Implementation details. We implemented the proposed framework in PyTorch on a workstation with two NVIDIA TITAN RTX GPUs (24 GB each). The VIFI-CLIP model (ViT-B/16 backbone) is fine-tuned in a few-shot setting with K=16 samples per class, using a batch size of 32 and a learning rate $2\times 10^{-6}$ for 50 epochs. We use the AdamW optimizer with a weight decay of 0.001 and apply cosine annealing for learning rate scheduling. Each video clip is represented by 32 uniformly sampled frames, which are resized to $224\times 224$. For the LLM, which classifies the student’s action sequence and classroom context as engaged or disengaged, we set the temperature to 0.1 to ensure a deterministic and non-creative output.
Our proposed framework consists of distinct tasks: action recognition, action parsing, and engagement classification. Each stage is evaluated with task-appropriate metrics. For action recognition, we report the top-1 accuracy. For action parsing, we adopt standard temporal action segmentation metrics, which fall into two categories: frame-based and segment-based metrics [11]. The frame-based metric, Mean over Frames (MoF), measures the ratio of correctly predicted frames to the total number of frames. However, the MoF metric suffers from two major problems. First, it is unreliable for imbalanced datasets, where long-duration action classes can dominate the sequence. Second, it fails to capture segment quality. A high MoF can still be achieved even if the action segments are fragmented (over-segmented). To account for the over-segmnetation phenomenon, we also report two segment-based metrics: Edit Score and $F1@\tau$. The Edit Score evaluates the similarity between the predicted and ground-truth action sequences without requiring exact frame alignment, capturing the correctness of action order. The $F1@\tau$ score measures the Intersection-over-Union (IoU) between predicted and ground-truth segments under threshold τ, counting a segment as a true positive only when its IoU exceeds $\tau$ (with $\tau\in\{10,25,50\}$). Multiple predictions overlapping a single ground-truth segment result in one true positive, and the rest are counted as false positives. Finally, we evaluate the engagement classification stage using precision, recall, and F1 score. These metrics are selected over simple accuracy to effectively assess performance in the presence of class imbalance between the engaged and disengaged labels.
4.2 Evaluation of the proposed framework stages
Stage #1: Action Recognition. We evaluated state-of-the-art video-language models, including XCLIP, TC-CLIP, and ViFi-CLIP on our student classroom action test set (containing 48 instances) under zero-shot and few-shot protocols. For the zero-shot setting, the ViFi-CLIP achieved 52.1% top-1 accuracy, which is far from satisfactory. We attribute this poor performance to the model’s fine-tuning on general web videos (Kinetics-400). While this bridges the image-video gap for general tasks, it fails to capture the subtle, fine-grained differences in our domain, such as differences between reading and writing. For the few-shot setting, as shown in Tab. 1, we note that the performance improves for all methods as the number of shots increases. Across most values of K, ViFi-CLIP achieves the highest accuracy, with particularly noticeable gains in the low- and mid-shot regimes, demonstrating that it can be effectively adapted to our specialized domain with minimal data. For example, ViFi-CLIP achieves gains of 2.1%, 4.2%, and 2.1%, compared to XCLIP at K=4, K=8, and K=16, respectively.
Table 1: Comparison of action recognition models.
Model Student Action Dataset
K=2 K=4 K=8 K=16
TC-CLIP [18] 43.8% 43.8% 77.1% 91.7%
XCLIP [22] 29.2% 72.9% 87.5% 95.8%
ViFi-CLIP [23] 39.6% 75.0% 91.7% 97.9%
Stage #2: Temporal action segmentation. Tab. 2 shows the comparison of Gemini 2.5 Pro, which is a multimodal large language model, and Gemini 2.5 flash with the proposed VLM action parsing on the student engagement dataset, which provides dense action labels for each 2-minute video. We curated a prompt describing the task that should be performed on the given 2-min video as shown in Fig. 8.
Figure 8: The prompt for temporal action segmentation.
It is noticeable from Tab. 2 that Gemini-2.5-pro achieves higher MoF, edit, and F1 scores compared to both Gemini-2.5-flash and the proposed VLM action parsing, while the proposed VLM action parsing outperforms Gemini-2.5-flash across all three metrics. Dense student action labeling is challenging due to fine-grained, visually similar actions (e.g., ‘looking at laptop’ vs. ‘looking down’) and cluttered environments. The task is further complicated by inter-student occlusion, where objects from one student’s space appear in another’s, creating ambiguous visual data. However, when we combined the visually similar actions e.g., writing on a notebook and reading, the performance is enhanced.
Table 2: Comparison of temporal action segmentation. Writing notes/reading means the two are merged into one class.
Method F1@10 F1@25 F1@50 Edit MoF
Gemini-2.5-pro 56.1 53.0 42.8 49.8 67.1
Gemini-2.5-flash 37.4 32.9 23.5 36.4 54.4
Our action parsing 44.6 39.3 27.1 42.0 58.9
Our action parsing (writing notes/Reading) 47.1 42.2 29.5 44.3 62.2
Stage #3: Engagement Classification.
To quantify the performance of the context-based engagement classifier, we first evaluate LLMs using the manual dense student action annotation, as shown in Tab. 3. We can notice that both Llama-3-8B and Gemma-2-9B achieve similar performance in terms of F1 score. The moderate performance of the LLMs on the disengaged class can be attributed to the inherent ambiguity of student engagement. In our dataset, some students were doodling on tablets rather than taking notes; experts labeled these cases as disengaged, but our action labels do not distinguish fine-grained behaviors such as writing versus drawing. Additionally, students who were daydreaming were sometimes annotated as listening because their posture suggested attention despite their lack of focus. Finally, student seated at the classroom periphery may appear to look forward without actually paying attention to the instructor. The context-aware engagement classifier achieves 92% F1 score when using ground truth action annotations. When employing automated action parsing, it attains 86% with Gemini-based parsing, Tab. 5, and 85% with the VLM-based parsing, Tab. 4. This performance gap is attributable to the difficulty of the action recognition task, as even the advanced Gemini-2.5-Pro model struggles to accurately parse the 2-minute-long student videos.
Table 3: Comparison of LLMs for context-based engagement classification based on manual annotation of student tubelets.
       LLM Recall Precision F1-score
Llama-3-8B disengaged 0.76 0.56 0.65
engaged 0.92 0.97 0.94
weighted avg 0.90 0.92 0.91
Gemma-2-9B disengaged 0.69 0.67 0.68
engaged 0.95 0.96 0.95
weighted avg 0.92 0.92 0.92
GPT-3.5 disengaged 0.71 0.52 0.60
engaged 0.91 0.96 0.93
weighted avg 0.89 0.90 0.89
Table 4: Comparison of LLMs for context-based engagement classification based on VLM action parsing of student tubelets.
       LLM Recall Precision F1-score
Llama-3-8B disengaged 0.55 0.39 0.45
engaged 0.88 0.93 0.91
weighted avg 0.84 0.87 0.85
Gemma-2-9B disengaged 0.78 0.36 0.49
engaged 0.81 0.96 0.88
weighted avg 0.80 0.89 0.83
GPT3.5 disengaged 0.76 0.32 0.45
engaged 0.78 0.96 0.86
weighted avg 0.78 0.88 0.81
Table 5: Comparison of LLMs for context-based engagement classification based on Gemini-2.5-Pro-based temporal action segmentation of student tubelets.
       LLM Recall Precision F1-score
Llama-3-8B disengaged 0.58 0.41 0.48
engaged 0.88 0.94 0.91
weighted avg 0.85 0.87 0.86
Gemma-2-9B disengaged 0.69 0.37 0.48
engaged 0.84 0.95 0.89
weighted avg 0.82 0.88 0.84
GPT3.5 disengaged 0.73 0.38 0.50
engaged 0.96 0.84 0.89
weighted avg 0.83 0.89 0.85
4.3 Ablation Studies
To ablate our model, we performed two experiments: a) we removed the classroom context to measure its contribution, and b) we compared our sequence-based method against a histogram-based counterpart. In this section, all engagement classification was performed by Llama-3-8B, using the ground-truth dense action labels for each 2-minute student video.
a) Engagement classification without classroom context: Tabs. 3 and 6 shows a comparison between the proposed peer-aware approach and its peer-free counterpart, respectively. It is noticeable that the proposed peer-aware approach outperforms its peer-free counterpart, achieving a 2% higher F1 score for the disengaged class. The prompt for engagement classification based on the student’s actions is shown in Fig. 10.
Figure 10: The prompt for engagement classification.
Table 6: Comparison of context-free engagement classifier based on manual annotation of student tubelets using Llama-3-8B.
       LLM Recall Precision F1-score
Llama-3-8B disengaged 0.75 0.51 0.61
engaged 0.90 0.96 0.93
weighted avg 0.88 0.91 0.89
Gemma-2-9B disengaged 0.73 0.60 0.66
engaged 0.93 0.96 0.95
weighted avg 0.91 0.92 0.91
GPT-3.5 disengaged 0.75 0.44 0.55
engaged 0.87 0.96 0.91
weighted avg 0.85 0.90 0.87
b) Input representation: The sequence-based and histogram-based input representations of a student’s time-stamped actions are compared in Tab. 7. We note that although the proposed sequence-based method’s overall F1 scores are comparable to that of its histogram-based counterpart, the sequence-based approach clearly outperforms the latter in terms of recall for the disengaged class (75% vs. 49%). This substantial difference supports our hypothesis that what really reflects student engagement is the temporal sequence of actions rather than an overall duration of an action represented by a histogram.
Table 7: Comparison of input representation for engagement classification based on manual annotation of student tubelets Using Llama-3-8B .
       Representation Recall Precision F1-score
Sequence disengaged 0.75 0.51 0.61
engaged 0.90 0.96 0.93
weighted avg 0.88 0.91 0.89
Histogram disengaged 0.49 0.77 0.60
engaged 0.98 0.93 0.96
weighted avg 0.92 0.91 0.91
5 Conclusion
We presented a novel context-aware framework to measure student behavioral engagement in a classroom setting. The proposed framework uses a Vision-Language Model (VLM) to parse student actions and a Large Language Model (LLM) to categorize engagement based on the resulting action sequence and the classroom context (i. e., peers’ activities). The experimental results demonstrated that context awareness is crucial to take into account because, depending on the context, the same action may indicate engagement or disengagement. On our created dataset, we evaluated the proposed framework and obtained F1 scores for engagement classification of 92%, 86%, and 85% with manual, Gemini-based, and modified VLM-based temporal action segmentation, respectively. Furthermore, the proposed framework overcomes the challenge of the scarcity of annotated data required to model the diversity of student behaviors.